### 23 Jan 2024 v1

# ClipSAM: CLIP and SAM Collaboration for Zero-Shot Anomaly Segmentation

Abstract
Recently, foundational models such as CLIP and SAM have shown promising performance for the task of Zero-Shot Anomaly Segmentation (ZSAS). However, either CLIP-based or SAM-based ZSAS methods still suffer from non-negligible key drawbacks: 1) CLIP primarily focuses on global feature alignment across different inputs, leading to imprecise segmentation of local anomalous parts; 2) SAM tends to generate numerous redundant masks without proper prompt constraints, resulting in complex post-processing requirements. In this work, we innovatively propose a CLIP and SAM collaboration framework called ClipSAM for ZSAS. The insight behind ClipSAM is to employ CLIP’s semantic understanding capability for anomaly localization and rough segmentation, which is further used as the prompt constraints for SAM to refine the anomaly segmentation results. In details, we introduce a crucial Unified Multi-scale Cross-modal Interaction (UMCI) module for interacting language with visual features at multiple scales of CLIP to reason anomaly positions. Then, we design a novel Multi-level Mask Refinement (MMR) module, which utilizes the positional information as multi-level prompts for SAM to acquire hierarchical levels of masks and merges them. Extensive experiments validate the effectiveness of our approach, achieving the optimal segmentation performance on the MVTec-AD and VisA datasets. Our code is public. [^2] https://github.com/Lszcoding/ClipSAM
1 Introduction
Zero-Shot Anomaly Segmentation (ZSAS) is a critical task in fields such as image analysis Fomalont (1999) and industrial quality inspection Mishra et al. (2021); Bergmann et al. (2022, 2019). Its objective is to accurately localize and segment anomalous regions within images, without relying on prior class-specific training samples. As a result, the diversity of industrial products and the uncertainty in anomaly types pose significant challenges for the ZSAS task.
Figure 1: Structural comparisons among different approaches for Zero-Shot Anomaly Segmentation. Top: CLIP-based approaches. Middle: SAM-based approaches. Bottom: Our ClipSAM approach that leverages the strengths of both CLIP and SAM methods.
With the emergence of foundational models such as CLIP Radford et al. (2021) and SAM Kirillov et al. (2023), the notable advancements have been achieved in Zero-Shot Anomaly Segmentation. As depicted in Figure 1, the CLIP-based approaches, like WinCLIP Jeong et al. (2023) and APRIL-GAN Chen et al. (2023a), determine the anomaly classification of each patch by comparing the similarity between image patch tokens and text tokens. While CLIP exhibits a strong semantic understanding capability, it is achieved by aligning the global features of language and vision, making it less suitable for fine-grained segmentation tasks Wang et al. (2022). Due to the fact that anomalies consistently manifest in specific regions of objects, the global semantic consistency inherent in CLIP is unable to achieve precise identification of the edges of local anomalies. On the other hand, researchers have explored the SAM-based approaches to assist the ZSAS task. SAM has superior segmentation capabilities and can accept diverse prompts, including points, boxes, and textual prompts, to guide the segmentation process. To this end, SAA Cao et al. (2023), as shown in Figure 1, utilizes the SAM with textual prompts to generate vast candidate masks and applies filters for post-processing. However, simple textual prompts may be insufficient for accurately describing anomalous regions, resulting in subpar anomaly localization performance and underutilization of SAM’s capabilities. Meanwhile, the ambiguous prompts lead to the generation of redundant masks, which requires further selection of correct masks.
Based on these observations, we innovatively propose a CLIP and SAM collaboration framework, first employing CLIP for anomaly localization and rough segmentation and then utilizing SAM and the localization information to refine the anomaly segmentation results. For the stage of CLIP, it is crucial to incorporate the fusion of language tokens and image patch tokens and model their dependencies to strengthen CLIP’s ability to segment anomaly parts, since cross-modal interaction and fusion have been proven to be beneficial for localization and object segmentation in several studies Jing et al. (2021); Xu et al. (2023); Feng et al. (2021). Further, several notable works Ding et al. (2019); Huang et al. (2019); Hou et al. (2020) have focused on enhancing the model’s local semantic comprehension by paying attention to row and column features. Motivated by these studies, we have developed a novel cross-modal interaction strategy that facilitates the interaction of text and visual features at both row-column and multi-scale levels, adequately enhancing CLIP’s capabilities for positioning and segmenting anomaly parts. For the stage of SAM, in order to fully harness its fine-grained segmentation capability, we exploit the localization abilities of CLIP to provide more explicit prompts in the form of both points and bounding boxes. This approach greatly enhances SAM’s ability to segment anomaly regions accurately. Further, we have noticed that SAM’s segmentation results often display masks with different levels of granularity, even when provided with the same prompt. To avoid the inefficient post-processing caused by further mask filtering, we propose a more efficient mask refinement strategy that seamlessly integrates different levels of masks, leading to enhanced anomaly segmentation results.
As a conclusion, we propose a novel two-stage framework named CLIP and SAM Collaboration (ClipSAM) for ZSAS. The structural comparisons with previous works are illustrated in Figure 1. In the first stage, we employ CLIP for localization and rough segmentation. To achieve the fusion of multi-modal features at different levels, we design the Unified Multi-scale Cross-modal Interaction (UMCI) module. UMCI aggregates image patch tokens from both horizontal and vertical directions and utilizes the corresponding row and column features to interact with language features to perceive local anomalies in different directions. UMCI also considers the interaction of language and multi-scale visual features. In the second stage, we exploit the CLIP’s localization information to guide SAM for segmentation refinement. Specifically, we propose the Multi-level Mask Refinement (MMR) module, which first extracts diverse point and bounding box prompts from the CLIP’s anomaly localization results and then uses these prompts to guide SAM to generate precise masks. Finally, we fuse these masks with the results obtained from CLIP based on different mask confidences. Our main contributions can be summarized as follows:
  • We propose a novel framework named CLIP and SAM Collaboration (ClipSAM) to fully leverage the characteristics of different large models for ZSAS. Specifically, we first use CLIP to locate and roughly segment the anomaly objects, and then refine the segmentation results with SAM and the positioning information.
  • To better assist CLIP in realizing desired localization and rough segmentation, we propose the Unified Multi-scale Cross-modal Interaction (UMCI) module, which learns local and global semantics about anomalous parts by interacting language features with visual features at both row-column and multi-scale levels.
  • To refine the segmentation results with SAM adequately, we designed the Multi-level Mask Refinement (MMR) module. It extracts point and bounding box prompts from the CLIP’s localization information to guide SAM in generating accurate masks, and fuse them with the results of CLIP to achieve fine-grained segmentation.
  • Extensive experiments on various datasets consistently validate that our approach can achieve new state-of-the-art zero-shot anomaly segmentation results. Particularly on the MVTec-AD dataset, our method outperforms the SAM-based method by +19.1↑ in pixel-level AUROC, +10.0↑ in $F_{1}$-max and +45.5↑ in Pro metrics.
2 Related Work
Figure 2: Overview of the proposed ClipSAM framework. ClipSAM includes two main processes: using CLIP for localization and rough segmentation, and using positioning information to prompt SAM to refine the segmentation results. These processes contain two important components: the Unified Multi-scale Cross-modal Interaction (UMCI) module and the Multi-level Mask Refinement (MMR) module. The UMCI module is employed for the interaction of language features with visual features of different directions and scales, facilitating CLIP’s ability to locate and segment anomaly objects. Meanwhile, the MMR module combines SAM, and uses point and box prompts extracted from location information to guide SAM to output the desired masks, and fuses them with the rough segmentation result obtained by CLIP.
2.1 Zero-shot Anomaly Segmentation
The methods for Zero-Shot Anomaly Segmentation (ZSAS) can be mainly divided into two categories. The first category is based on CLIP. As the pioneering work, WinCLIP Jeong et al. (2023) calculates the similarity between image patch tokens and textual features for ZSAS. Further, APRIL-GAN Chen et al. (2023a) employs linear layers to better align features of different modalities. AnoVL Deng et al. (2023) and ANOMALYCLIP Zhou et al. (2023) propose to enhance the generalization of text. SDP Chen et al. (2023b) proposes to address noise in the encoding process of the CLIP image encoder. The second category is based on SAM. Specifically, SAA Cao et al. (2023) utilizes text prompts for SAM to generate vast candidate masks and uses a complex evaluation mechanism to filter out irrelevant masks.
However, relying solely on CLIP or SAM may lead to certain limitations. For instance, CLIP-based methods struggle with precise segmentation of local anomalies, while SAM-based methods heavily rely on specific prompts. To address these drawbacks, we explore the collaboration mechanism of CLIP and SAM and propose a novel framework to leverage their strengths. Besides, we further design the Unified Multi-scale Cross-modal Interaction (UMCI) module and Multi-level Mask Refinement module to better exploit the specific ability of CLIP and SAM, respectively.
2.2 Foundation Models
Recently, there has been an increasing focus and attention on foundation models. Various foundation models have achieved satisfactory performance on kinds of downstream tasks Devlin et al. (2018); Brown et al. (2020). Notably, CLIP Radford et al. (2021) and SAM Kirillov et al. (2023) have emerged as two representative models with impressive zero-shot reasoning capabilities in classification and segmentation tasks, respectively. More specifically, CLIP focuses on aligning multi-modal features and possesses robust semantic understanding abilities for both language and vision, while SAM excels in achieving fine-grained segmentation based on different prompts. Recently, Yue et al. (2023) attempts to establish a connection between SAM’s image encoder and CLIP’s text encoder for surgical instrument segmentation, Wang et al. (2023) attempts to merge SAM and CLIP to facilitate downstream tasks. These works highlight the critical importance of exploring collaboration among foundational models.
2.3 Cross-modal Interaction
In the field of multi-modal learning, cross-modal interaction is becoming increasingly important. Specifically, Hu et al. (2016) concatenates features from different modalities and uses convolutions for multi-modal information fusion. Further, STEP Chen et al. (2019) establishes correlations between important areas in the image and relevant keywords in the text to enhance the fusion of cross-modal information. BRINet Feng et al. (2021) exchanges cross-modal information between different blocks of the encoder to facilitate image segmentation. The success of cross-modal interaction in diverse domains has motivated us to explore it in the context of zero-shot anomaly segmentation. To effectively address the challenge of localizing abnormal regions within an object, we introduce the Unified Multi-scale Cross-modal Interaction module, taking into account the interaction between text and both row-column and multi-scaled visual features.
3 Methodology
3.1 CLIP and SAM Collaboration
CLIP has a strong semantic understanding of different modalities, while SAM can easily detect edges of fine-grained objects, both of which are important for anomaly segmentation. In this paper, we present a novel CLIP and SAM collaboration framework called ClipSAM, which aims to boost the performance of ZSAS. The overview architecture is illustrated in Fig. 2. Specifically, we leverage the CLIP for initial rough segmentation and utilize it as a constraint to refine the segmentation results with SAM. In Sec. 3.2, we introduce the Unified Multi-scale Cross-modal Interaction (UMCI) module within the CLIP stage to achieve accurate rough segmentation and anomaly localization. In Sec. 3.3, we design the Multi-level Mask Refinement (MMR) module, which incorporates the guidance from CLIP to facilitate SAM in generating more precise masks for achieving fine-grained segmentation. In addition, the optimization function of the overall framework is discussed in Sec. 3.4.
3.2 Unified Multi-scale Cross-modal Interaction
In our ClipSAM framework, the CLIP encoder is employed to process both text and image inputs. For a specific pair of text and image, the encoder generates two outputs: $L\in\mathbb{R}^{C_{t}\times 2}$ and $P_{i}\in\mathbb{R}^{H\times W\times C}$. Here, $L$ represents the textual feature vector, in line with WinCLIP Jeong et al. (2023), reflecting two categories of normal and abnormal. $P_{i}$ denotes the patch tokens derived from the i-th stage of the encoder. For more details of the textual feature $L$, please refer to Appendix A.
As discussed in Sec. 1, cross-modal fusion has been proven to be beneficial for object segmentation. To achieve this, the Unified Multi-scale Cross-modal Interaction (UMCI) module is designed. Specifically, the UMCI model consists of two parallel paths: the ${Strip}$ ${Path}$ and the ${Scale}$ ${Path}$. The ${Strip}$ ${Path}$ captures both row- and column-level features of the patch tokens to precisely pinpoint the location. The ${Scale}$ ${Path}$ focuses on grasping the image’s global features of various scales, enabling a comprehensive understanding of the anomaly. More details are described below.
(1) ${Strip}$ ${Path}$. Denote the inputs of a specific UMCI module as textual feature vector $L$ and patch tokens ${P}$. We first process the patch tokens to grasp the visual features. The image features are projected to align with the text features in dimension, resulting in $\hat{P}\in\mathbb{R}^{H\times W\times C_{t}}$. To extract row- and column-level features from $\hat{P}$, we apply two average pooling layers, with kernel sizes of ${1\times W}$ and ${H\times 1}$ respectively:
$v_{row}=conv_{1\times 3}(Avg\_Pool_{1\times W}(\hat{P})),$ (1)
$v_{col}=conv_{3\times 1}(Avg\_Pool_{H\times 1}(\hat{P})),$
where $v_{row}\in\mathbb{R}^{H\times c_{h}}$ and $v_{col}\in\mathbb{R}^{W\times c_{h}}$ are the row-level and column-level features. ${H}$ and ${W}$ denote the height of the vertical feature and the width of the horizontal feature.
We then focus on the internal process of the ${Strip}$ ${Path}$. Take $v_{row}$ for example, we apply convolution layers to the text features $L$, obtaining $t_{row}^{1},t_{row}^{2}\in\mathbb{R}^{c_{h}\times 2}$. $t_{row}^{1}$ and $t_{row}^{2}$ serve as the text feature input of the subsequent Scaled Dot-Product Attention mechanism Vaswani et al. (2017), a crucial component for the interaction between the language and visual domains. Specifically, we implement a two-step attention mechanism to efficiently predict the language perception of the pixels in $v_{row}$ (normal or abnormal), denoted as $M_{row}\mathbb{R}^{H\times c_{h}}$. More details are provided in Appendix B. In parallel, $v_{col}$ is similarly processed to obtain $M_{col}\in\mathbb{R}^{W\times c_{h}}$.
We then utilize the bilinear interpolation to expand $M_{row}$ and $M_{col}$ to their original scales and combine the results:
$M_{row,col}=conv_{3\times 3}(B(M_{row})+B(M_{col})),$ (2)
where ${B}$ symbolizes the bilinear interpolation layer. This process results in $M_{row,col}\in\mathbb{R}^{H\times W\times c_{h}}$.
(2) ${Scale}$ ${Path}$. In this path, the image features are also projected to $\hat{P}\in\mathbb{R}^{H\times W\times C_{t}}$. Then we apply two average pooling layers with kernel sizes of $s_{1}$ and $s_{2}$ to grasp the visual features of different scales:
$v_{{g}_{1}}=conv_{3\times 3}^{{g}_{1}}(Avg\_Pool_{{s}_{1}\times{s}_{1}}(\hat{P})),$ (3)
$v_{{g}_{2}}=conv_{3\times 3}^{{g}_{2}}(Avg\_Pool_{{s}_{2}\times{s}_{2}}(\hat{P})),$
where $v_{{g}{1}}\in\mathbb{R}^{h{{g}{1}}\times w{{g}{1}}\times c{h}}$ and $v_{{g}{2}}\in\mathbb{R}^{h{{g}{2}}\times w{{g}{2}}\times c{h}}$ represent visual features at different scales.
For the internal process of ${Scale}$ ${Path}$, we consider $v_{{g}_{1}}$ for example. The text features are processed by convolution layers and yield $t_{{g}_{1}}^{k},t_{{g}_{1}}^{v}\in\mathbb{R}^{c_{{g}_{1}}\times 2}$. We then obtain the language perception of the pixels $M_{{g}_{1}}\in\mathbb{R}^{h_{{g}_{1}}\times w_{{g}_{1}}\times c_{h}}$ by the attention with $v_{{g}_{1}}$ as the query, $t_{{g}_{1}}^{k}$ as the key and $t_{{g}_{1}}^{v}$ as the value. More details are provided in Appendix B. Similar to the ${Strip}$ ${Path}$, we utilize the bilinear interpolation to resize and combine $M_{{g}_{1}}$ and $M_{{g}_{2}}$:
$M_{{g}_{1},{g}_{2}}=conv_{3\times 3}^{{g}_{1},{g}_{2}}(B(M_{{g}_{1}})+B(M_{{g}_{2}})),$ (4)
(3) ${Dual-path}$ ${Fusion}$. After the ${Strip}$ ${Path}$ and ${Scale}$ ${Path}$, we have obtained the pixel-wise predictions $M_{row,col}$ and $M_{{g}_{1},{g}_{2}}$, providing comprehensive location and semantics information about the anomaly. The last step of the UMCI module involves fusing these results to get the rough segmentation of the anomalous region. Specifically, we introduce a residual connection from the input patch token $\hat{P}$, and fuse it with the pixel-wise predictions by a convolution layer:
$v_{ori}=conv_{3\times 3}^{ori}(\hat{P}),$ (5)
$M_{all}=conv_{3\times 3}^{all}(concat(v_{ori},M_{row,col},M_{{g}_{1},{g}_{2}})).$
We employ a Multi-Layer Perceptron as the segmentation head, and the rough segmentation of the anomalous regions is mathematically described as:
$O=MLP(ReLU(M_{all}+\hat{P})).$ (6)
where $O\in\mathbb{R}^{H\times W\times 2}$ denotes the segmentation output of a specific UMCI module, and dimension 2 represents the classification of the foreground anomalous parts and the background. Assume there are $n$ stages in the encoder, and denote $O_{i}$ as the segmentation output of stage ${i}$. Then the final segmentation results can be calculated as $O=\frac{1}{n}\textstyle\sum_{i=1}^{n}O_{i}$.
3.3 Multi-level Mask Refinement
Figure 3: The results produced by SAM with different spatial prompts. As we can see, constraining SAM with the spatial prompt that represents points and boxes as a whole leads to better results.
With the rough segmentation $O$ from the CLIP phase, we propose the Multi-level Mask Refinement (MMR) module to extract point and box prompts to guide SAM to generate accurate masks. In the MMR module, the foreground of the rough segmentation, denoted as $O_{f}\in\mathbb{R}^{H\times W}$, is firstly post-processed with a binarization step to obtain a binary mask ${O_{b}(x,y)}$. Denote $v(x,y)_{f},x\in{H},y\in{W}$ as the value of a specific pixel in $O_{f}$, then the value of each pixel $v(x,y)_{b}$ in ${O_{b}(x,y)}$ can be calculated as:
$v(x,y)_{b}=\begin{cases}&1,\text{if }v(x,y)_{f}>threshold\\ &0,otherwise\end{cases}$ (7)
where $threshold$ represents the binary threshold, and the value 1 corresponds to the anomalous pixels. Within the connected areas of this binary mask, we identify some boxes and points to provide spatial prompts for SAM. For point selection, $m$ random points are chosen, represented as $S_{p}=[(x_{p_{1}},y_{p_{1}}),\dots,(x_{p_{m}},y_{p_{m}})]$, where $(x_{p_{i}},y_{p_{i}})$ represents the position of the i-th point. Boxes are generated based on the size of connected regions in the binary mask, with the i-th box denoted by $S_{b_{i}}=[(x_{b_{i}},y_{b_{i}},h_{b_{i}},w_{b_{i}})]$. The complete set with $q$ boxes is represented as $S_{b}=[S_{b_{1}},\dots,S_{b_{q}}]$.
With the point prompts $S_{p}$ and box prompts $S_{b}$, we explore the optimal prompt sets for SAM. As can be seen in Figure 3, employing either points or boxes alone leads to biased results. In contrast, the combined application of both points and boxes can yield more precise and detailed segmentation. Therefore, in our ClipSAM framework, we use $S=S_{b}\cup{S_{p}}$ as the prompt set for SAM. With the original image $I$ and spatial prompts $S$ as inputs, SAM generates encoded features ${z}_{i}$ and ${z}_{s}$. The decoder within SAM then outputs the refined masks and corresponding confidence scores:
$({masks},{scores})=\mathbb{D}^{sam}({z}_{i}\mid{z}_{s}).$ (8)
Each box shares the same point constraints, resulting in $q$ distinct segmentation masks. In our ClipSAM framework, SAM is configured to produce three masks with varying confidence scores for each box, represented as $masks=[(m_{1}^{1},m_{1}^{2},m_{1}^{3});\dots;(m_{q}^{1},m_{q}^{2},m_{q}^{3})]$ and $scores=[(s_{1}^{1},s_{1}^{2},s_{1}^{3});\dots;(s_{q}^{1},s_{q}^{2},s_{q}^{3})]$. The final fine-grained segmentation result $O_{final}$ is obtained by normalizing the fusion of rough segmentation and the refined masks:
$O_{final}=Norm(O+\sum_{i=1}^{q}\sum_{j=1}^{3}m_{i}^{j}\times s_{i}^{j}).$ (9)
3.4 Objective Function
In our ClipSAM framework, the only part involving training is the UMCI module. To effectively optimize this module, we employ the Focal Loss Lin et al. (2017) and the Dice Loss Milletari et al. (2016), both of which are well-suited for segmentation tasks.
Focal Loss. Focal loss is primarily applied to address class imbalance problem, a common challenge in segmentation tasks. It is appropriate for anomaly segmentation because, usually, the anomaly merely occupies a small fraction of the entire object. The expression of Focal loss is:
$l_{focal}=-\frac{1}{H\times W}\textstyle\sum_{i=0}^{H\times W}(1-p_{i})^{\gamma}\log{(p_{i})},$ (10)
where $p_{i}$ is the predicted probability for a pixel being abnormal, and $\gamma$ is a tunable parameter and set to 2 in our paper.
Dice Loss. Dice Loss calculates a score based on the overlap between the target area and the model’s output. This metric is also effective for class imbalance issue. Dice Loss can be calculated as:
$l_{dice}=1-\frac{1}{N}\frac{2\times{\textstyle\sum_{i=1}^{N}}y_{i}\hat{y}_{i}}{{\textstyle\sum_{i=1}^{N}}y_{i}^{2}+{\textstyle\sum_{i=1}^{N}}\hat{y}_{i}^{2}},$ (11)
where $N=H\times W$ is the total number of pixels in features.
Total Loss. We set separate loss weights for each stage and the total loss can be expressed as:
$l_{all}=\textstyle\sum_{i=1}^{n}\lambda_{i}(l_{focal}^{i}+l_{dice}^{i}),$ (12)
where $i$ denotes the index of stages, $\lambda_{i}$ is the loss weight of the i-th stage. The CLIP encoder in our implementation consists of 4 stages in total, and we set the loss weights for these stages at $0.1$, $0.1$, $0.1$, and $0.7$ respectively.
4 Experiments
MVTec-AD VisA
Base model Method AUROC $F_{1}$-max AP PRO AUROC $F_{1}$-max AP PRO
CLIP-based Approaches WinCLIP 85.1 31.7 - 64.6 79.6 14.8 - 56.8
APRIL-GAN 87.6 43.3 40.8 44.0 94.2 32.3 25.7 86.8
SDP 88.7 35.3 28.5 79.1 84.1 16.0 9.6 63.4
SDP+ 91.2 41.9 39.4 85.6 94.8 26.5 20.3 85.3
SAM-based Approaches SAA 67.7 23.8 15.2 31.9 83.7 12.8 5.5 41.9
SAA+ 73.2 37.8 28.8 42.8 74.0 27.1 22.4 36.8
CLIP & SAM ClipSAM(Ours) 92.3 47.8 45.9 88.3 95.6 33.1 26.0 87.5
Table 1: Performance comparison of different kinds of ZSAS approaches on the MVTec-AD and VisA datasets. Evaluation metrics include AUROC, $F_{1}$-max, AP, and PRO. Bold indicates the best results.
4.1 Experimental Setup
Datasets. In this study, we conduct experiments on two commonly-used datasets of industrial anomaly detection, namely VisA Zou et al. (2022) and MVTec-AD Bergmann et al. (2019), which encompass a diverse range of industrial objects categorized as normal or abnormal. We follow the same training setup as existing zero-shot anomaly segmentation studies Jeong et al. (2023); Chen et al. (2023a) to evaluate the performance of our method. Specifically, the model is first trained on the MVTec-AD dataset and then tested on the VisA dataset, and vice versa. Additional experimental results on other datasets are provided in Appendix C.
Metrics. Following Jeong et al. (2023), we employ widely-used metrics, i.e., AUROC, AP, $F_{1}$-max, and PRO Bergmann et al. (2020), to provide a fair and comprehensive comparison with existing ZSAS methods. Specifically, AUROC reflects the model’s ability to distinguish between classes at various threshold levels. AP quantifies the model’s accuracy across different levels of recall. $F_{1}$-max is the harmonic mean of precision and recall at the optimal threshold, implying the accuracy and coverage of the model. PRO assesses the proportion of correctly predicted pixels within each connected anomalous region, offering insights into the model’s local prediction accuracy. Higher values of these metrics mean better performance of the evaluated method.
Implementation details. In the experiments, the pre-trained ViT-L-14-336 model released by OpenAI, which consists of 24 Transformer layers, is utilized for CLIP encoders. We extracted the image patch tokens after each stage of the image encoder (i.e., layers 6, 12,18, and 24) for the training of our proposed UMCI module, respectively. The optimization process is conducted on a single NVIDIA 3090 GPU using AdamW optimizer with the learning rate of ${1\times 10^{-4}}$ and the batch size of 8 for 6 epochs. For SAM, we use the ViT-H pre-trained model.
Figure 4: Comparison of visualization results among ClipSAM, CLIP-based, and SAM-based methods on the MVTec-AD dataset. Our ClipSAM performs much better on the location and boundary of the anomaly segmentation.
Figure 5: Visualization of the results of each step of our ClipSAM collaboration framework. ClipSAM first uses CLIP for rough segmentation and then uses SAM for refinement.
4.2 Experiments on MVTec-AD and VisA
Comparison with state-of-the-art approaches. In this section, we evaluate the effectiveness of our proposed ClipSAM framework for ZSAS on the MVTec-AD and VisA datasets. Table 1 shows the comprehensive comparison between our proposed ClipSAM and the state-of-the-art ZSAS methods Jeong et al. (2023); Chen et al. (2023a); Cao et al. (2023); Chen et al. (2023b) on different datasets and various metrics. It can be concluded that our proposed ClipSAM outperforms existing state-of-the-art methods in all four metrics. Taking the MVTec-AD dataset as an example, our proposed ClipSAM outperforms the advanced CLIP-based method SDP+ by 1.1%, 5.9%, 6.5% and 2.7% on the AUROC, $F_{1}$-max, AP and PRO metrics respectively. Compared to the SAM-based approach, our method exhibits superior performance benefits, i.e., improvements of 19.1%, 10.0%, 17.1%, and 45.5% for the metrics. On the VisA dataset, our proposed method similarly shows an overall performance enhancement, demonstrating the effectiveness and generalization of our ClipSAM.
Qualitative comparisons. We provide some visualization of ZSAS results in Figure 4 to further demonstrate the effectiveness of the proposed method. For comparison, we also show the segmentation visualization of APRIL-GAN (CLIP-based method) and SAA (SAM-based method). It can be observed that APRIL-GAN can roughly locate the anomalies but fails to provide excellent segmentation results. In contrast, SAA can perform the segmentation well but cannot cover the anomalous region accurately. Compared with these methods, our proposed ClipSAM provides accurate localization as well as good segmentation results. More visualization results are provided in Appendix D. To better understand the role of each module in the ClipSAM framework, we also visualize the rough segmentation results of the CLIP phase and the processed prompts fed into the SAM phase in Figure 5. It shows that CLIP performs a rough segmentation of abnormal parts and generates corresponding prompts based on their locations to complete SAM’s further refinement of the results. Please refer to Appendix E for more details.
Components of ClipSAM AUROC $F_{1}$-max AP PRO
UMCI only w/strip 90.8 44.3 34.9 79.6
only w/scale 90.9 44.4 42.7 81.9
Module w/o UMCI 60.4 19.7 25.0 34.7
w/o MMR 91.8 46.7 44.7 84.8
ClipSAM(Ours) 92.3 47.8 45.9 88.3
Table 2: Ablation study of different components in our ClipSAM framework on MVTec-AD dataset. Bold indicates the best results.
Hyperparameters AUROC $F_{1}$-max AP PRO
Hidden dim ($c_{h}$) 194 91.8 45.6 43.8 83.1
256 91.7 46.4 44.6 83.6
384 92.3 47.8 45.9 88.3
Kernel size ($s_{1}$ $s_{2}$) 2 & 4 91.9 45.7 45.3 85.1
3 & 9 92.3 47.8 45.9 88.3
6 & 10 91.8 46.8 44.7 84.9
Threshold ($thr$) 0.45 91.5 43.3 44.9 82.5
0.47 92.3 47.8 45.9 88.3
0.50 91.7 44.6 43.4 83.2
Table 3: Ablation study of different hyperparameters used in our ClipSAM framework on MVTec-AD dataset. Bold indicates the best results in the UMCI module. $c_{h}$ represents the hidden dimension of the convolutional layer. $s_{i}$ denotes the kernel size of the average pooling layer used in the scale path. $thr$ means the threshold for binarization.
Figure 6: Visualization of anomaly localization and rough segmentation by CLIP with the UMCI module and CLIP with similarity calculation.
4.3 Ablation Studies
In this section, we conduct several ablation studies on the MVTec-AD dataset to further explore the effect of different components and the experiment settings on the results in the proposed ClipSAM framework.
Effect of components. Table 2 shows the results of the ablation study of different components in ClipSAM. Specifically, we first explore the impact of preserving only the strip path or the scale path in the UMCI module. Subsequently, the performance of removing either the UMCI or MMR module from the framework is tested. It can be found that removing a path in the UMCI module can lead to performance degradation, and removing the scale path has a greater impact, reflecting the necessity of combining two paths in the UMCI module. At the module level, removing the MMR module will slightly lower performance. Comparing this result with SDP+, we can surprisingly find that the rough segmentation results of the CLIP phase yield even better performance than CLIP-based methods. Figure 6 shows the visualization comparison between the rough segmentation and APRIL-GAN (Since SDP is not open source). The first two columns of Figure 6 indicates that UMCI can locate the anomalies more accurately, and the last two columns shows that UMCI provides better segmentation. In comparison with the MMR module, removing the UMCI module means regarding the similarity-based segmentation as the rough segmentation result. However, as shown in Figure 4, the similarity-based segmentation cannot provide text-aligned patch tokens and accurate local spatial prompts for the subsequent MMR module. This results in a performance collapse, which demonstrates the important role of the UMCI module in the ClipSAM framework.
Effect of hyperparameters. We explore the effects of various hyperparameters on our ClipSAM framework and record the results in Table 3. An analysis of the roles of each hyperparameter in the ClipSAM framework is provided based on the results. Hidden Dimension (${c_{h}}$) determines the output feature dimension of the convolution layer. Larger ${c_{h}}$ values contribute to the effective interpretation of visual features by the model. Kernel size (${s_{i}}$) affects the size of multi-scale visual features, which should be moderate to provide easy-to-understand visual context. Threshold Value ($thr$) primarily impacts the initial binarized segmentation of the SAM phase. The setting of $thr$ should also be moderate: a small value may cause the non-anomalous regions to be misclassified and thus unable to generate accurate masks for SAM; a large value may cause some anomalies to be ignored and not detected.
5 Conclusion
We propose the CLIP and SAM Collaboration (ClipSAM) framework to solve zero-shot anomaly segmentation for the first time. To cascade the two foundation models effectively, we introduce two modules. One is the UMCI module, which explicitly reasons anomaly locations and achieves rough segmentation. The other is the MMR module, which refines the rough segmentation results by employing the SAM with precise spatial prompts. Sufficient experiments showcase that ClipSAM provides a new direction for improving ZSAS by leveraging the characteristics of different foundation models. In future work, we will further investigate how to integrate knowledge from different models to enhance the performance of zero-shot anomaly segmentation.